OpenStudy (anonymous):

pi((x+10)^2-(x+6)^2) simplifying

OpenStudy (inkyvoyd):

LOLOL pi(x^2+20x+100-x^2-12x-36)=pi(20x-12x+100-36)=pi(8x+64)

OpenStudy (anonymous):

lol what

OpenStudy (inkyvoyd):

pi(8x+64)